OpenStudy (anonymous):

Are 3x + 33 / x + 11 and 3 equivalent? Using complete sentences explain why or why not.

OpenStudy (lyrae):

\[(3x + 33) / (x + 11) = 3(x + 11) / (x + 11) = 3 \]

OpenStudy (anonymous):

so they are equivalent?

OpenStudy (anonymous):

ok thank u